OpenStudy (anonymous):

F(3)+2...all it says is "Compute the following functional expressions" how do I do that?

OpenStudy (callisto):

Does the question give you what F(3) is or what F(x) is equal to ?

OpenStudy (anonymous):

hint is: Given that f(x) = 3x2 + 2x, find f(h). Everywhere that my formula has an "x", I now plug in an "h": f(x) = 3x2 + 2x f( ) = 3( )2 + 2( ) f(h) = 3(h)2 + 2(h) = 3h2 + 2h Given that f(x) = 3x2 + 2x, find f(x + h). Everywhere that my formula has an "x", I now plug in an "x + h". f(x) = 3x2 + 2x f( ) = 3( )2 + 2( ) f(x + h) = 3(x + h)2 + 2(x + h) = 3(x2 + 2xh + h2) + 2x + 2h = 3x2 + 6xh + 3h2 + 2x + 2h